jcolonna07:

Calculate the average daily balance based on the following information: 29-day billing cycle 4/17 Billing date and previous balance $1,100 4/27 Payment $700 4/29 Charge $300 5/7 Payment $50

mhanifa:

You need to use the formula: Average daily balance = Sum of (day#*balance) / number of days 1. Day 1 is 4/17 with balance of $1100 until 4/26, this gives us 10 days and the sum of: $1100*10 = $11000 2. Day 11 is 4/27, the balance changes: $1100 - $700 = $400 Until next balance change we have 2 days and the sum is: $400*2 = $800 3. Day 13 is 4/29, the balance changes: $400 + $300 = $700 Until next balance change on 5/7 we have 8 days and the sum is: $700*8 = $5600 4. Day 21 is 5/7, the balance change: $700 - $50 = $650 This stays for the rest of period, thus 9 days, and the sum is: $650*9 = $5850 Add up all the sums above and divide by 29 to get your answer.